OpenStudy (anonymous):

i am stumpted on a question , about what an equation would be for an x y chart any body who thinks they can help?

OpenStudy (saifoo.khan):

is that related to Linear Law?

OpenStudy (anonymous):

im not quite sure its new to me , i could give you the numbers for the chart?

OpenStudy (oaktree):

Are you talking about equations of the form y=mx+b?

OpenStudy (anonymous):

yes

OpenStudy (oaktree):

OK, give me the numbers.

OpenStudy (saifoo.khan):

if you have a chart, you can post a photo here or use the draw tool to draw.

OpenStudy (oaktree):

Just give me two points on the line, or a point and the slope.

OpenStudy (anonymous):

x, 0, 1, 2, 3, 4, 5, 6, 7, 8 y, 0, 3, 12, 33, 72, 135, 228, 375, 528

OpenStudy (oaktree):

Wow. OK, I only needed two.

OpenStudy (oaktree):

Let's start with finding the slope, or the "m" part of y=mx+b

OpenStudy (oaktree):

So, m=\[(y _{2}-y_{1})/(x_{2}-x_{1})\]

OpenStudy (anonymous):

how did you get that tho

OpenStudy (oaktree):

Where two points are \[(x_{2},y_{2}) and (x_{1}, y_{1})\]

OpenStudy (oaktree):

Oh, that's just a theorem.

OpenStudy (oaktree):

Good so far?

OpenStudy (anonymous):

not at all im not suposed to graph i am just supposed to guess using the chart

OpenStudy (oaktree):

No, we're not graphing. We're solving for the slope and the y-intercept, so that we'll have an equation like y=5x-3.

OpenStudy (oaktree):

OK?

OpenStudy (oaktree):

Well, thanks for leaving in the middle. The answer is y=(1/5)x+(14/5).